OpenStudy (anonymous):

the lenght of a base of a trapezoid are 12cm and 26cm find the lenght of a median?

OpenStudy (anonymous):

\[\text{Length Of Median} = \frac{1}{2} \times (\text{Sum Of Length Of Bases)}\]

OpenStudy (anonymous):

@anna14 , what are the lengths of bases you are given??

OpenStudy (anonymous):

@anna14 interested in replying??

OpenStudy (anonymous):

26 and 12 cm

OpenStudy (anonymous):

If you will reply, then in two minutes your question will take in solving..

OpenStudy (anonymous):

can you answer my question..please

OpenStudy (anonymous):

Yes, you are right.. Now what is their sum, can you add the two lengths you are given with..??

OpenStudy (anonymous):

I cannot give you your answer, but I can tell you how to reach at your answer..

OpenStudy (anonymous):

Hey man, come on, add the two lengths you said to me..

OpenStudy (anonymous):

ahyy okey i get it thanks..:)

OpenStudy (anonymous):

Really? So, what you got as your answer??

OpenStudy (anonymous):

See, don't mind, If you lie, you will die...

OpenStudy (anonymous):

76cm

OpenStudy (anonymous):

How you got 76 cm??

OpenStudy (anonymous):

@hba stay out of it, ha ha ha..

OpenStudy (anonymous):

You are stopping a man from learning something..

OpenStudy (anonymous):

hey..f.y.i im not a boy ..im a gurll..are u blind @waterineyes heheh

OpenStudy (anonymous):

so..:P

OpenStudy (anonymous):

Ha ha ha,. That is why I am a teacher.. @anna14 , hba is making me loose all my concentration.. Okay, sorry for that, you are a girl and even nice girl if you add 12 + 26 correctly..

OpenStudy (anonymous):

Tell me : \[\huge \color{red}{12 + 26 = ?}\]

OpenStudy (anonymous):

38

OpenStudy (anonymous):

Yep well done.. Now can you half it, Divide it by 2..

OpenStudy (anonymous):

So I think what anna meant was 12+26=38 * 2 =76 or something?

OpenStudy (hba):



OpenStudy (anonymous):

\[\huge \color{green}{ \frac{38}{2} = ?}\]

OpenStudy (anonymous):

76

OpenStudy (anonymous):

you do not insult ever thought I turned analyzed

OpenStudy (anonymous):

This is where you are doing mistake..

OpenStudy (anonymous):

Have I insulted you??

OpenStudy (anonymous):

38 * 1/2 = one half of 38

OpenStudy (anonymous):

Sorry, if you felt bad..

OpenStudy (anonymous):

@hba: you turning me like a child.sorryim only a person..:(

OpenStudy (anonymous):

@hba:im just a person..

OpenStudy (anonymous):

@hba:im just a person..

OpenStudy (anonymous):

@hba:im just a person..

OpenStudy (anonymous):

@hba get lost ..

OpenStudy (anonymous):

@anna14 ignore hba..

OpenStudy (anonymous):

hba chill

OpenStudy (hba):

@anna14 Please ignore me

OpenStudy (anonymous):

@hba this is not fair..

OpenStudy (anonymous):

@hba:have you studied?

OpenStudy (hba):

@waterineyes Everything is fair :P

OpenStudy (hba):

@anna14 I am uneducated.